HUBBLE SPACE TELESCOPE DAILY REPORT #1002 PERIOD COVERED: 0800 EST Daily Status Report as of 300/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 5041 (GHRS Sensitivity Monitor - Cycle 3) The High Resolution Spectrograph was used to make 24 ultraviolet observations of standard star AGK81D266 to monitor the sensitivity of the HRS, which is expected to change slowly with time. The onboard target acquisition was successful. All 24 observations were received and displayed in the OSS. They were reported to look normal. 1.2 Completed FOS/RD/BL 5120 (HST Imagery and Spectroscopy of the Dusty SMC H II Regions N88A) The Faint Object spectrograph was used on both the Red and Blue sides to make observations of the dusty hydrogen II regions of N88A which is located in the Large Magellanic Cloud. It is hoped that these observations will help learn more of the conditions within this nebula and the properties of the dust which it contains. Some of the observations still remain on the science tape recorder, and will be analyzed later. Where possible, the OSS monitored the other observations, and assessed the data quality. No problems were reported. 1.3 Completed Two Sets of FOS/BL 4110 (The Physics and Massive O-Stars Different Parent Galaxies) The Faint Object Spectrograph (Blue Side) was used to make observations of giant O class stars in galaxy NGC-346-6. Ultraviolet observations from 1150 A0 to 2300 A0 were made. The binary target acquisition was successful. The following peakup found the maximum count in the seventh group. It is possible that the target was not well centered in the slit during the observations. A recentering event was also reported to have occurred during subsequent observations. No other anomalies were reported. 1.4 Completed WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.6 Completed PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) The Planetary Camera was used to make internal flat observations to check for quantum efficiency hysteresis. The observations were executed as schedule, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 0 Successful : 0 2.3 Significant HSTARs: None 2.4 Operations Note: The SMGT #14(A) test was initiated yesterday, Tuesday Oct. 26, and will end on Thursday, Oct. 28, 1993. All operations are proceeding normally, and no anomalies have been reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/BL 4110 (The Physics of Massive O-Stars in Different Parent Galaxies) 3.2 Complete PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.3 Complete FOC/96 5045 (FOC Observations to Monitor Desorption - Cycle 3) 3.4 Complete FOC/96 5029 (FOC Plate Scale - Cycle 3) 3.5 Complete FOC/96 4969 (Search for QSOs Suitable for Subsequent Observation of He II) 3.6 Complete HSP/POL 4888 (Remnant Stars In Supernova Remnants - Aug. - Cycle 3 High) 3.7 Complete WFC 5076 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 2) 3.8 Complete HSP/POL 5039 (GHRS Dark Noise Monitoring - Cycle 3) 4.9 Complete HSP/POL 4889 (Gravitational Lenes Part 1 - Aug. - Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-